A 55.0 kg students is in an elevator and pushes the button to go up. Briefly the acceleration is 0.300 g as it starts moving up. If she is standing on a scale what would it read while this acceleration is happening?
Answer #1

When the acceleration is upward with acceleration, a = 0.3 g

Scale reading = M*g + M*a

= M*(g+a)

= M*(g+0.3g)

=M*1.3*g

= 55*1.3*9.8

=700.7 N

Answer: 700.7 N